Required: Based upon what you find, answer the following questions: A. For Discount Goods, do liabilities or shareholders eqDiscount Goods 2021 Balance Sheet Store (AU) Big Store ✓ Discount Goods 42,308 78,225 Account name Total current liabiliti..Big Store 2021 Balance Sheet Store (All) Big Store Discount Goods 45,996 51,928 Account name Total current liabiliti.. Total
Answer #1

Balance Sheet of Discount Goods 2021

Particiulars Amount Particulars Amount
Shareholders' Equity 69130 Noncurrent Assets 106571

Noncurrent Liabilities

73358 Current Assets 78225
Current Liabilites 42308
Total 184796 184796

Balance Sheet of Big Store 2021

Particiulars Amount Particulars Amount
Shareholders' Equity 66296 Noncurrent Assets 113905

Noncurrent Liabilities

41677 Current Assets 45996
Current Liabilites 51928
Total 159901 159901

A. For Discount Goods, liabilities provide the greater proportion of the company's Assets

B. For Big Store, liabilities provide the greater prportion of Company's Assets

C. For Current liabilities to total liabilities,

For Discount Goods Ratio = Current Liabilities / Total Liabilities

                                       = 42308 / (73358 + 42308)

                                      = 42308/115666

                                      = 0.3658

                                      = 37%

For Big Store Ratio = Current Liabilities / Total Liabilities

                             = 51928 / (41677 + 51928)

                              = 51928/93605

                              = 0.5548

                             = 55%

For Current liabilities to total liabilities, Big Store reported the highest ratio of 55 %

D. For Current Assets to Current Liabilities

For Discount Goods Ratio = Current Assets / Current Liabilites

                                      = 78225/42308

                                       = 1.85

                                       = 185%

For Big Store Ratio = Current Assets / Current Liabilites

                                      = 45996 / 51928

                                      = 0.8858

                                      = 89%

For Current Assets to Current liabilities, Discount Goods reported the highest ratio of 185%